# T-FAKE: Synthesizing Thermal Images for Facial Landmarking

### v1

@@ 3.2 The Dataset
Having trained $T_{\theta}$, we use it to thermalize the FAKE dataset which results in the thermal dataset T-FAKE. For each thermal condition, the dataset contains 100k images featuring 2k unique subjects. Each image is available in a simulated ‘cold’ and a ‘warm’ environment adding up to 200k images. This number of frames and subjects surpasses the numbers in other thermal landmarking datasets [29, 54, 12, 43, 32, 33, 3] known to the authors. In contrast to thermal landmarking datasets with a fixed distance between the subjects and the camera [29, 43, 33], the distance varies within an estimated range of 80cm to 150cm for the T-FAKE dataset. Moreover, the synthesized images are based on facial scans with a wide range of age and ethnicity and a balanced number of people identifying as male or female [52]. In addition to the original ground truth sparse landmark annotations, we add dense annotations by transferring Mediapipe [35] predictions for the FAKE to the T-FAKE dataset. This makes our dataset the first densely annotated thermal landmarking dataset. While we found many of the large thermal landmarking datasets to be non-public [54, 12, 43], our complete dataset with both thermal conditions and annotations is available for download. In combination with the original FAKE dataset, we can train multimodal facial landmarkers jointly on RGB and thermal images as described in the next section. Table 1 displays a detailed comparison between thermal landmarking datasets
@@ 4 Facial Landmarking
Learning with synthetic images has proven to be very useful in facial landmarking [54, 52, 53]. Based on the proposed T-FAKE and the original FAKE dataset, we can thus train a supervised thermal or multimodal landmark prediction model for thermal and RGB images using synthetic data only. Moreover, we can adapt our model to new annotation conventions by training a small-scale label adaptation model. The data flow for these two models is visualized in Fig. 2.
@@ 4.1 Landmark Prediction
$-\log p(y|\mu,\sigma)=\log\left(2\pi\sigma^{2}\right)+\frac{\|\mu-y\|_{2}^{2}}{2\sigma^{2}},$
@@ 4.2 Label Adaptation
A common approach that is especially crucial for synthetic trained data is label adaptation [52, 53]. It allows us to adapt to different landmark conventions between the training dataset and an evaluation dataset. Given two sequences $(\hat{\mathbf{y}}_{k},\mathbf{y}_{k})_{k=1}^{K}$ with predicted landmarks $\hat{\mathbf{y}}_{k}\coloneqq(\hat{y}^{l}_{k})_{l=1}^{\hat{L}}\in\mathbb{R}^{2\hat{L}}$, and with ground truth landmarks $\mathbf{y}_{k}:=(y^{l}_{k})_{l=1}^{L}\in\mathbb{R}^{2\hat{L}}$ with possibly $\hat{L}\neq L$, we aim to train a third label adaptation model $T_{\zeta}:\mathbb{R}^{2\hat{L}}\to\mathbb{R}^{2{L}}$. Here, we use the loss
@@ 5.1 Implementation Details
Thermalization. We train on the RGB+Thermal SEJONG dataset recorded in lab conditions and the large-scale synthetic FAKE dataset featuring in-the-wild RGB images. We train on a resolution of $256\times 256$. For the patch-based regularizers, we use a patch size of $8$. We disregard the background by excluding synthetic patches based on the ground truth background segmentation and completely black real patches. For our multi-scale approach, we sum the regularizer over $5$ scales with a downsampling factor of $0.5$. Furthermore, facial temperatures depend on the surrounding temperature.
Thus, we train a ‘warm’ and a ‘cold’ model with different reference temperature values for the segmentation-based regularizers, see Fig. 3. For data augmentation, we use the same random rotations and cropping for the natural RGB and thermal images. Moreover, we apply random color changes, blurring, and shadow augmentations [38] exclusively to the natural RGB images. Next, we apply random rotations and cropping for the synthetic RGB images. Here, we also fill holes in the original ‘glasses’ segmentation masks showing outlines of the frame only to highlight transparent, but heat-blocking glass or plastic, see Fig. 3 (bottom right). Lastly, we replace the original background with a black background based on the known semantic segmentation. We apply our two models to the FAKE dataset at the original $512\times 512$ resolution to generate the T-FAKE dataset. This results in the up-to-date largest thermal landmarking dataset with 200k images of 2k unique subjects and two thermal settings, see Table 1.
Landmarker. We simultaneously train our landmark prediction model using a sparse 70-point and dense 468-point landmark convention with an image size of $224\times 224$ on our new T-FAKE thermal dataset and the original FAKE RGB dataset. Additionally, we include random landmark positions on a texture dataset [16] as negative examples into the dataset to increase the learned uncertainty $\sigma^{2}$ on images without faces. Due to the limited, public availability of the dense ground truth in the FAKE dataset [52], we use the 468-point predictions of the MediaPipe model for the synthetic RGB images as our dense ground truth dense annotations. We make use of spatial augmentations, including random shear, rotations, resizing, and cropping to allow the landmarker to learn a large variation of face orientations without the need for a dedicated face detector. Furthermore, we employ Gaussian blur, color jitter, and global pixel offset for thermal images during training. On thermal images, we fill in the background with random textures from the texture dataset [16]. During inference, we use a multi-scale sliding window evaluation to generalize our model to varying image sizes. We downsample iteratively with a factor of $0.75$ until the height or the width is smaller than $224$. For each image scale, we run our model on sliding windows of size $224\times 224$ with a stride of $20$. For our final landmark prediction, we pool all predictions and use the landmark with the smallest predicted standard deviation across all scales and all sliding windows.
@@
Label Adaptation. Our label adaptation network is trained solely on the two-dimensional landmark predictions and ground truth annotations of a training split of a benchmark dataset and not on the images. This dataset is later excluded from the final evaluation. We do not use the predicted standard deviation at this stage. Again, we use random rotations and shear transformations during training.
@@
Evaluation Metrics for Landmarking. We report the normalized mean error (NME) [52, 48, 56] for evaluation, i.e., the mean absolute error (MAE) normalized with an image-dependent constant to account for varying image sizes. Often, normalization is performed by the distance between the outermost eye landmarks (NME Interocular) [52, 48]. However, the CHARLOTTE benchmark dataset contains many examples with side profile views where only one of the eyes is visible. Therefore, we use the average height and width of the bounding box around the face for normalization [56] (NME W/H). Apart from the NME, we report the failure rates for each of the reference methods. The failure rate is the percentage of frames where no faces could be found. Frames without recognized faces are not included nor weighted in the NMEs. The method of rejection varies for each landmarker. We use vanilla parameters except for MediaPipe [35], where we set the minimal detection confidence to $0.01$. For our method, we evaluate it i) on all frames without failure rate and ii) on all frames where the average uncertainty scores $\sigma_{l}$ of the predicted landmarks $l$, $1\leq l\leq L$, is less than $\bar{\sigma}=6\times 10^{-4}$.
@@
Metric Method Charlotte High Charlotte Low Charlotte Side Charlotte Front Charlotte Full
NME Aachen [28] 0.1677 0.2405 0.2942 0.1731 0.2054
3FabRec [5] 0.2862 0.3195 0.1810 0.4259 0.3033
3FabRec* 0.2646 0.3272 0.1921 0.4015 0.2967
Mediapipe [17] 0.1900 0.2745 0.2879 0.1865 0.2262
Mediapipe* 0.1400 0.2621 0.1838 0.2164 0.2023
FA [6] 0.1070 0.2790 0.1222 0.2508 0.1875
FA* 0.1007 0.2207 0.1075 0.2123 0.1593
TFW [32] 0.0787 0.1249 0.0701 0.1311 0.1012
SF-TL54 U-Net [33] 0.2476 0.3134 0.1757 0.3596 0.2829
SF-TL54 Dlib [33] 0.2159 0.2174 0.1365 0.2741 0.2167
Ours dense ($\bar{\sigma}<\infty$) 0.0805 0.1285 0.0846 0.1258 0.1050
Ours ($\bar{\sigma}<\infty$) 0.0740 0.1346 0.0684 0.1420 0.1051
Ours ($\bar{\sigma}<6\times 10^{-4}$) 0.0729 0.1139 0.0628 0.1219 0.0923
Failure rate (%) Aachen [29] 49.32 48.72 72.85 25.15 49.01
Mediapipe [35] 25.05 46.85 50.09 22.40 36.25
Mediapipe* 12.80 14.07 24.97 1.92 13.45
FA [6] 12.95 27.43 21.71 19.07 20.39
FA* 0.83 10.45 4.75 6.80 5.77
TFW [32] 13.14 21.55 19.26 15.66 17.46
SF-TL54 U-NET / Dlib [33] 29.29 22.57 38.16 13.50 25.84
Ours ($\bar{\sigma}<6\times 10^{-4}$) 1.63 16.45 9.15 9.34 9.25
@@
Table 2: Results on the CHARLOTTE dataset. Pre-processing with the full pre-processing stack for RGB landmarkers is indicated by *. The low resolution (3434 frames) and high resolution (3250 frames) are defined by the median image width of 200 pixels. We denote the confidence threshold with $\bar{\sigma}$. Note that we include all predictions for $\bar{\sigma}<\infty$ and 3FabRec resulting in a failure rate of 0 by design.
@@
Figure 4: Results on examples from the CHARLOTTE dataset [3] of FA and Mediapipe with the full pre-processing stack and our models without label adaptation. While the RGB methods can handle some challenging poses, they produce broken landmarks on certain difficult images.
@@ 5.2 Benchmark Datasets for Landmarking
To compare thermal performance, we use the CHARLOTTE dataset due to its high variability in resolution, pose, and background temperature. The dataset is annotated with either 72 landmarks for frontal images or 43 landmarks for lateral images. The 43-point landmarks represent the visible landmarks in profile pictures with our information on orientation. The dataset contains a total of 7685 annotated frames with different resolutions with a minimal one of around 70x60. For evaluation, we split the CHARLOTTE dataset based on image size and perspective. The CHARLOTTE High condition contains all images larger than the median image width of 200 pixels and CHARLOTTE Low all smaller images. The condition CHARLOTTE Side includes all images with profile annotations (43 landmarks), and CHARLOTTE Front contains all frontal faces (73 landmarks). In addition, we evaluate on the established RGB landmark dataset 300W [48] to verify RGB landmark prediction performance. It uses the classical 68-point convention. Here, we crop images to the pre-defined bounding boxes around each face for our evaluation. The average image resolution is about 290 pixels.
300W
Method NME W/H NME Interocular
3FabRec [5] 0.0376 0.0646
FA [6] 0.0618 0.1083
Ours 0.0330 0.0577
Ablation Study
GB 0.0303 0.0523
Table 3: Comparison of RGB performance of our method. We trained the label adaptation on 300 random samples of the combined W300 outdoor and indoor scenes (600 images) and evaluated the remaining images.
Regularization FID
$\lambda_{R}=0,\lambda_{W}=0$ 0.3325 $\pm$ 0.0039
$\lambda_{R}=0,\lambda_{W}=0.01C$ 0.3139 $\pm$ 0.0035
$\lambda_{R}=1,\lambda_{W}=0$ 0.2719 $\pm$ 0.0054
$\lambda_{R}=1,\lambda_{W}=0.01C$ 0.2666 $\pm$ 0.0037
@@
Table 4: Influence of the regularizers on the thermalization. The parameter $\lambda_{W}$ corresponds to the impact of the patch-based regularizer (Eq. (2)) and $\lambda_{R}$ corresponds to the segmentation-based regularizer (Eq. (3)). We use $C=(5\cdot 8^{2})^{(-1)}$.
Metric Method Charlotte High Charlotte Low Charlotte Side Charlotte Front Charlotte Full
NME RGB 0.1055 0.2675 0.1241 0.2534 0.1887
RGB* 0.0933 0.2682 0.1312 0.2348 0.1824
Thermal Only 0.0832 0.1334 0.0677 0.1503 0.1090
Table 5: Ablation results on the CHARLOTTE dataset. Pre-processing with the full pre-processing stack for RGB landmarkers is indicated by *. Here, we do not exclude low-confidence images. Hence, we report the NME on all images without failure rate
5.3 Comparison to State-of-the-Art
@@
We compare the performance of our landmarker trained on our full synthetic datasets against the state-of-the-art models on the CHARLOTTE dataset. Here, we include not only relevant thermal landmark detection systems but also RGB landmark detection systems. A previous study has shown that the performance of RGB landmarkers on thermal images can be significantly boosted by applying a stack of unlearned linear filters before prediction [12]. Consequently, the application of this pre-processing stack of linear filters allows us to include RGB landmarkers in our benchmark and to draw a fair comparison beyond task-specific thermal landmarkers. Because the CHARLOTTE dataset uses its own landmark convention, we use label adaptation for all methods for the image evaluation. We train each label adaptation model with a 1000-image training split on CHARLOTTE similar to [52] and exclude the training split from the final evaluation. Note that this approach allows us to include our sparse model and our dense model in the evaluation. To benchmark the RGB performance of the multimodal model, we evaluate our sparse model predictions against state-of-the-art RGB sparse landmarkers on the 300W dataset.
@@
CHARLOTTE performance. The NME performance as well as the failure rates of the different methods varies drastically for the different CHARLOTTE splits. We observe the lowest performance across the landmarkers on the Front condition, followed by the low-resolution images. The FA and Mediapipe RGB methods show a performance that is comparable with the SF-TL54 and Aachen thermal landmarkers. Using a pre-processing stack [12], they consistently outperform them on the High condition in terms of NME and on Full, Front, Low and Side in terms of failure rate. Our model achieves state-of-the-art performance for sparse thermal landmarking across all different conditions. The only model with comparable performance is TFW which, however, only predicts a set of 5 landmarks. While the NME of TFW is slightly lower than our method for $\bar{\sigma}<\infty$ for Low, Front, and Full, our method outperforms TFW on all dataset splits in terms of NME and failure rate for $\bar{\sigma}<6\times 10^{-4}$. These results are presented in Table 2 and examples are visualized in Fig. 4.
300W performance. Additionally, we evaluate the RGB models on the 300W dataset in Table  3. Our model, trained on a combination of thermal and RGB facial images, demonstrates competitive performance on RGB images, being outperformed only by the original model trained exclusively on RGB data. The overall performances are consistent with the literature [52].

### v2

@@ 3.2 The Dataset
Having trained $T_{\theta}$, we use it to thermalize the FAKE dataset which results in the thermal dataset T-FAKE. For each thermal condition, the dataset contains 100k images featuring at least 2k unique subjects. Each image is available in a simulated ‘cold’ and a ‘warm’ environment adding up to 200k images. This number of frames and subjects surpasses the numbers in other thermal landmarking datasets [34, 63, 13, 52, 37, 38, 3] known to the authors. In contrast to thermal landmarking datasets with a fixed distance between the subjects and the camera [34, 52, 38], the distance varies within an estimated range of 80cm to 150cm for the T-FAKE dataset. Moreover, the synthesized images are based on facial scans with a wide range of age and ethnicity and a balanced number of people identifying as male or female [61]. In addition to the original ground truth sparse 3DA-2D landmark annotations, we add dense 2D annotations by transferring Mediapipe [40] predictions for the FAKE to the T-FAKE dataset. This makes our dataset the first densely annotated thermal landmarking dataset. While we found many large thermal landmarking datasets to be non-public [63, 13, 52], our complete dataset with both thermal conditions and annotations is available for download.
In combination with the original FAKE dataset, we can train multimodal facial landmarkers jointly on RGB and thermal images as described in the next section. Table 1 displays a comparison between thermal landmarking datasets.
Figure 4: Challenging frames from the FAKE dataset [61] before (first row) and after thermalization without any regularization (second row) and our domain-adaptive model (third row). Thermalized backgrounds are removed based on given segmentation. The last row displays the segmentation masks with gray levels which are proportional to the reference temperature values used in regularizer (3). The first three images exemplify the ‘cold’ setup with colder noses and the last two the ‘warm’ setup with warmer noses.
@@ 4 Facial Landmarking
Learning with synthetic images has proven to be very useful in facial landmarking [63, 61, 62]. Based on the proposed T-FAKE and the original FAKE dataset, we can train a supervised thermal or multimodal landmark prediction model for thermal and RGB images using synthetic data only. Moreover, we adapt a model to new annotation conventions by training a small-scale label adaptation model. The data flow for these models is visualized in Fig. 3.
@@ 4.1 Landmark Prediction
$-\log p(y|\mu,\sigma)=\log\left(2\pi\sigma^{2}\right)+\frac{\|\mu-y\|_{2}^{2}}{2\sigma^{2}},$ (9)
@@ 4.2 Label Adaptation
A common approach that is especially crucial for synthetic trained data and comparison of 2D and 3DA-2D landmarks is label adaptation [61, 62]. It allows us to adapt to different landmark conventions between the training dataset and an evaluation dataset. Given two sequences $(\hat{\mathbf{y}}_{k},\mathbf{y}_{k})_{k=1}^{K}$ with predicted landmarks $\hat{\mathbf{y}}_{k}\coloneqq(\hat{y}^{l}_{k})_{l=1}^{\hat{L}}\in\mathbb{R}^{2\hat{L}}$, and with ground truth landmarks $\mathbf{y}_{k}:=(y^{l}_{k})_{l=1}^{L}\in\mathbb{R}^{2\hat{L}}$ with possibly $\hat{L}\neq L$, we aim to train a third label adaptation model $T_{\zeta}:\mathbb{R}^{2\hat{L}}\to\mathbb{R}^{2{L}}$. Here, we use the loss
@@ 5.1 Implementation Details
Thermalization. We train on the RGB+Thermal SEJONG dataset recorded in lab conditions and the large-scale synthetic FAKE dataset featuring in-the-wild RGB images. We train on a resolution of $256\times 256$.
The resolution-agnostic U-net allows to apply our two models trained on $256\times 256$ images to the FAKE dataset at the original $512\times 512$ resolution to generate the T-FAKE dataset. This results in the up-to-date largest thermal landmarking dataset with 200k images of at least 2k unique subjects and two thermal settings, see Table 1. We refer to the supplementary for details.
Landmarker. We simultaneously train our landmark prediction model using a sparse 70-point 3DA-2D and dense 478-point 2D landmark convention with an image size of $224\times 224$ on our new T-FAKE thermal dataset and the original FAKE RGB dataset. Due to the limited, public availability of the dense 3DA-2D ground truth in the FAKE dataset [61], we use the 478-point 2D predictions of the MediaPipe model for the synthetic RGB images as our dense ground truth dense annotations.
During inference, we use a multi-scale sliding window evaluation to generalize our model to varying image sizes without direct face tracking. For our final landmark prediction, we pool all predictions and use the landmark with the smallest predicted standard deviation across all scales and all sliding windows. We refer to the supplementary for more details.
@@
Label Adaptation. Our label adaptation network is trained solely on the two-dimensional landmark predictions and ground truth annotations of our random training split of a benchmark dataset and not on the images. This dataset is later excluded from the final evaluation. We do not use the predicted standard deviation at this stage. Again, we use random rotations and shear transformations during training.
@@
Evaluation Metrics for Landmarking. We report the normalized mean error (NME) [61, 58, 68] for evaluation, i.e., the mean absolute error (MAE) normalized with an image-dependent constant to account for varying image sizes. Often, normalization is performed by the distance between the outermost eye landmarks (NME IO) [61, 58]. However, the CHARLOTTE benchmark dataset contains many examples with side profile views where only one of the eyes is visible. Therefore, we use the average height and width of the bounding box around the face for normalization [68] (NME W/H). Apart from the NME, we report the failure rates for each of the reference methods. The failure rate is the percentage of frames where no faces could be found. Frames without recognized faces are not included nor weighted in the NMEs. The method of rejection varies for each landmarker. We use vanilla parameters except for MediaPipe [40], where we set the minimal detection confidence to $0.01$. For our method, we evaluate it i) on all frames without failure rate and ii) on all frames where the average uncertainty scores $\sigma_{l}$ of the predicted landmarks $l$, $1\leq l\leq L$, is less than a fixed value $\bar{\sigma}$.
@@
Metric Method Training Dataset High Low Side Front Full
NME W/H $\downarrow$ 3FabRec [5] 300W 0.0931 0.1084 0.0675 0.1344 0.1009
3FabRec* 300W 0.0814 0.1558 0.0791 0.1602 0.1196
Star [65] 300W 0.0707 0.1021 0.0673 0.1064 0.0868
Star* 300W 0.0639 0.1273 0.0703 0.1227 0.0965
Mediapipe [18] undisclosed 0.1900 0.2745 0.2879 0.1865 0.2262
Mediapipe* undisclosed 0.1400 0.2621 0.1838 0.2164 0.2023
FA [6] LS3D-W 0.1070 0.2790 0.1222 0.2508 0.1875
FA* LS3D-W 0.1007 0.2207 0.1075 0.2123 0.1593
FAKE[61, 62] FAKE 0.1055 0.2675 0.1241 0.2534 0.1887
FAKE* FAKE 0.0933 0.2682 0.1312 0.2348 0.1824
Aachen [33] AACHEN 0.1677 0.2405 0.2942 0.1731 0.2054
TFW [37] TFW 0.0787 0.1249 0.0701 0.1311 0.1012
SF-TL54 U-Net [38] SF-TL54 0.2476 0.3134 0.1757 0.3596 0.2829
SF-TL54 Dlib [38] SF-TL54 0.2159 0.2174 0.1365 0.2741 0.2167
Ours dense ($\bar{\sigma}<\infty$) FAKE + T-FAKE 0.0854 0.1395 0.0764 0.1500 0.1132
Ours ($\bar{\sigma}<\infty$) FAKE + T-FAKE 0.0683 0.1239 0.0645 0.1294 0.0969
Ours ($\bar{\sigma}<6\times 10^{-4}$) FAKE + T-FAKE 0.0671 0.1118 0.0593 0.1179 0.0889
Ours ($\bar{\sigma}<2.2\times 10^{-4}$) FAKE + T-FAKE 0.0661 0.1018 0.0551 0.1059 0.0815
Failure rate (%) $\downarrow$ Aachen [34] AACHEN 49.32 48.72 72.85 25.15 49.01
Mediapipe [40] undisclosed 25.05 46.85 50.09 22.40 36.25
Mediapipe* undisclosed 12.80 14.07 24.97 1.92 13.45
FA [6] LS3D-W 12.95 27.43 21.71 19.07 20.39
FA* LS3D-W 0.83 10.45 4.75 6.80 5.77
TFW [37] TFW 13.14 21.55 19.26 15.66 17.46
SF-TL54 U-NET / Dlib [38] SF-TL54 29.29 22.57 38.16 13.50 25.84
Ours ($\bar{\sigma}<6\times 10^{-4}$) FAKE + T-FAKE 0.98 10.34 6.94 4.64 5.79
Ours ($\bar{\sigma}<2.2\times 10^{-4}$) FAKE + T-FAKE 6.52 32.76 23.33 16.68 20.00
@@
Table 2: Results on CHARLOTTE dataset splits. Pre-processing with the pre-processing stack for RGB landmarkers is indicated by *. Best NME value is in bold. We denote the confidence threshold with $\bar{\sigma}$. We include all predictions for $\bar{\sigma}<\infty$, Star and 3FabRec resulting in an optimal failure rate of 0 by design. Methods in gray used additional supervision from bounding boxes generated with the ground truth landmarks. RGB datasets are indicated by .
@@
Figure 5: CHARLOTTE [3] examples using FA and Mediapipe with pre-processing stack and our models without label adaptation. While RGB methods can handle some challenging poses, they may produce broken landmarks on difficult images.
@@ 5.2 Benchmark Datasets for Landmarking
To compare thermal performance, we use the CHARLOTTE dataset due to its high variability in resolution, pose, and background temperature. The dataset is annotated with either 72 2D landmarks for frontal images or 43 2D landmarks for lateral images. The 43-point landmarks represent the visible landmarks in profile pictures with our information on orientation. We use a total of 7684 annotated frames with different resolutions with a minimal one of around 70x60 from the dataset. For evaluation, we propose to split the CHARLOTTE dataset based on image size and perspective. The High condition contains all images larger than the median image width of 200 pixels and Low all smaller images. The condition Side includes all images with profile annotations (43 landmarks), and Front contains all frontal faces (73 landmarks). In addition, we evaluate on the established RGB landmark dataset 300W [58] to verify RGB landmark prediction performance. It uses the classical 68-point convention. Here, we crop images to the pre-defined bounding boxes around each face for our evaluation. We follow the protocol from [65] for 300W evaluation. The average image resolution is about 290 pixels.
300W (NME IO $\downarrow$)
Method Full Comm. Chal.
Mediapipe [40] 0.0562 0.0469 0.0947
FA [6] 0.0435 0.0380 0.0661
Star [65] 0.0287 0.0252 0.0430
Ours 0.0522 0.0468 0.0743
Table 3: Comparison of RGB performance of our method. We trained the label adaptation on the 300W train split and evaluated each method on the 300W test split [24, 65]. For Star, we use the authors’ model pre-trained on 300W. NME has been normalized with respect to interocular distance [61, 58]. Methods with label adaptation are put in italic.
Model FID $\downarrow$ TUFTS MSE [48] $\downarrow$ DRIVE-IN MSE $\downarrow$ [13]
$\lambda_{R}=0,\lambda_{W}=0$5028 .0378 .1117
$\lambda_{R}=0,\lambda_{W}=0.01C$3146 .0360 .1021
$\lambda_{R}=1,\lambda_{W}=0$1706 .0357 .0545
$\lambda_{R}=1,\lambda_{W}=0.01C$1598 .0353 .0580
Pix2Pix [26] .2003 .0421 .1453
CycleGAN [66] .2228 .0521 .0730
CUT [49] .2140 .0468 .1370
QS-Attn [23] .1173 .0409 .1303
@@
Table 4: Influence of the regularizers on the thermalization and comparison to other models. The parameter $\lambda_{W}$ corresponds to the impact of the patch-based regularizer (Eq. (2)) and $\lambda_{R}$ corresponds to the segmentation-based regularizer (Eq. (3)). We use $C=(5\cdot 8^{2})^{(-1)}$. The best value is in bold and the second-best is underlined.
5.3 Comparison for Landmarking
@@
We compare the landmarker performance with training on our full synthetic datasets against state-of-the-art pre-trained models on the CHARLOTTE dataset. Besides relevant thermal landmarkers we include RGB landmarkers. Performance of RGB landmarkers on thermal images can be boosted by applying a stack of unlearned linear filters before prediction [13]. Consequently, application of this pre-processing stack allows us to include RGB landmarkers in our benchmark and to draw a fair comparison beyond task-specific thermal landmarkers. We use label adaptation for all methods to account for different landmark conventions for the image evaluation. We train each label adaptation model with a 1000-image training split on CHARLOTTE similar to [61] and exclude the training split from the final evaluation. Note that this approach allows us to include our sparse model and our dense model in the evaluation. To benchmark the RGB performance of the multimodal model, we evaluate our sparse model predictions against state-of-the-art RGB sparse landmarkers on the 300W dataset.
@@
CHARLOTTE performance. The NME performance as well as the failure rates of the different methods varies drastically for the different CHARLOTTE splits. We list the dataset on which each of the methods has been trained in Table 2. We observe the lowest performance across the landmarkers on the Front condition, followed by the low-resolution images. The FA and Mediapipe RGB methods show a performance that is comparable with the SF-TL54 and Aachen thermal landmarkers. Using a pre-processing stack [13], they consistently outperform them on the High condition in terms of NME and on Full, Front, Low and Side in terms of failure rate. Our model achieves state-of-the-art performance for sparse thermal landmarking across all different conditions. The only model with comparable performance is TFW which, however, only predicts a set of 5 landmarks. While the NME of TFW is slightly lower than our method for $\bar{\sigma}<\infty$ for Low, Front, and Full, our method outperforms TFW on all dataset splits in terms of NME and failure rate for $\bar{\sigma}<6\times 10^{-4}$. These results are presented in Table 2 and examples are visualized in Fig. 5. The 300W-trained RGB methods Star and 3FabRec are close to our method and TFW, however, they have been supplied with the ground truth landmarks for tight bounding box computation. This gives an advantage for label adaptation of broken profile images as well as evaluation on broken and shifted ground truth in CHARLOTTE.
300W performance. Additionally, we evaluate the performance of our multimodal model on the 300W dataset in Table  3. Our model, trained on a combination of thermal and RGB facial images with label adaptation performs between MediaPipe and FA with label adaptation and all methods are outperformed by Star. It is important to note, that the bounding box has only been optimized for Star and our model has been trained without centered faces and a large variability of head positions in the input frames.
5.4 Comparison for Thermalization
Our approach aims to overcome the limitations of thermal datasets to fixed camera angles, controlled lighting, and stable temperatures. However, this data limitation hinders the evaluation of ‘in-the-wild’ images. Therefore, we report three different metrics. Firstly, we report the average Fréchet Inception Distance (FID) [21] between all available thermal SEJONG images and five different equally-sized thermalized FAKE subsets. The FID measures image translation without ground truth image pairs.Here, it compares the perceptual similarity of the thermal SEJONG dataset and a dataset of thermalized FAKE images. Additionally, we report the mean-squared error (MSE) for two paired thermal datasets (with filtered background) to evaluate actual out-of-distribution prediction quality. Here, we use TUFTS [48] and 1000 randomly selected images from DRIVE-IN [13]. TUFTS has been recorded under conditions similar to SEJONG and DRIVE-IN features side views from people in cars. Table 4 compares our final model ($\lambda_{R}=1$, $\lambda_{W}=0.01C$) to two supervised models trained with SEJONG only, a baseline implementation ($\lambda_{R}=0$, $\lambda_{W}=0$) and a Pix2Pix model [26], and three unsupervised models trained with SEJONG and FAKE, namely CycleGAN [66], CUT [49] and QS-Attn [23]. The MSE shows that our model outperforms all other models for temperature prediction and displays domain adaptation for the out-of-distribution DRIVE-IN data recorded outside a lab. Moreover, the resulting T-FAKE images display better perceptual similarity than images produced by all other methods except for QS-Attn. However, QS-Attn fails at temperature prediction, especially for the non-frontal DRIVE-IN images.